OpenStudy (anonymous):

When solving the system x + 2y = 9 and x - y = 6, which expression could be substituted for x in the first equation? Answer 9 - x y - 9/2 9 - 2y 1/2 y + 9

OpenStudy (anonymous):

Its 9-2y

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

Because you subtract y from both sides to get X alone.

OpenStudy (anonymous):

okay thanks! what about this one Solve the system. What is the value of x? y = 3x and 4x + 2y = 30 Answer 3 -3 9 -9

OpenStudy (anonymous):

- y + 4x + 2y = - 3x + 30 y = 30 - 7x 30 - 7x = 3x 10x = 30 x = 3 So your answer is 3.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Your welcome :)